OpenStudy (anonymous):

Match each equation with its equivalent equation. Match Term Definition six x equals twenty‾four A) x divided by twelve equals three x divided by six equals six B) four x equals sixteen x divided by two equals sixteen C) x divided by eight equals four eight x equals sixteen D) four x equals eight

OpenStudy (anonymous):

@SACAPUNTAS

OpenStudy (anonymous):

Is this off a paper or a screenshot or something? I know what you're supposed to do but the way you've got it written it out here hurts my brain.

OpenStudy (anonymous):

6X=24 A.)X/12=3 X/6=6 B.)4X=16 X/2=16 C.)X/8=4 8X=16 D.)4X=8

OpenStudy (anonymous):

Much better! :D

OpenStudy (anonymous):

All you really have to do is find x for each value on the left and then find the equivalent value of x on the right. For example, \(6x = 24\) We can divide both sides by 4 and find out that \(x = 4\) So which one the right has \(x = 4\)? \(x / 12 = 3\)? No, \(x = 12\times 3 = 36\). (Remember that for later) \(4x = 16\), yes, x = 4. So the first one matches with B. I think you know what to do now? :)

OpenStudy (anonymous):

Oops, I got a bit ahead of myself. I meant divide both sides by 6. 24/6 = 4.

OpenStudy (anonymous):

YEAH THANKS!

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

\[6x = 24\]\[\frac{6x}{6} = \frac{24}{6}\]\[x = 4\] It was a typo when I said divide by 4.

OpenStudy (anonymous):

i really dont get this can you do it for me please? i have tons of other work

OpenStudy (anonymous):

@SACAPUNTAS

OpenStudy (anonymous):

No, I'm not going to just do it for you. I can explain how but that's all I'll do.

OpenStudy (anonymous):

what about we do it together?

OpenStudy (anonymous):

Ok. Can you solve \(\frac{x}{6} = 6\)?